OpenStudy (anonymous):

Which of the following statements are true regarding functions? Check all that apply. A. The vertical line test may be used to determine whether a function is one-to-one. B. A sequence is a function whose domain is the set of real numbers. C. A function is a relation in which each value of the input variable is paired with exactly one value of the output variable. D. The horizontal line test may be used to determine whether a function is one-to-one. I know A is true, I'm not positive about B or C, and I know D is false.

OpenStudy (crashonce):

C isnt always true for example, when you plug one value into a quadratic for y, it may provide two solutions so C isnt always true

OpenStudy (anonymous):

so it's just A and B? @CrashOnce

OpenStudy (crashonce):

not sure about B though @ikram002p a little help please

OpenStudy (zzr0ck3r):

A is not true

OpenStudy (zzr0ck3r):

the vertical line test is to make sure its a function we use the horizontal line test to see if its 1-1

OpenStudy (zzr0ck3r):

the definition of one to one is that f(a)=f(b) imples a=b. If a line fails the horizontal line test then what that means there are at least 2 different x values mapping to the same y value.

OpenStudy (zzr0ck3r):

i.e. f(a) = f(b) but a \(\ne\) b

OpenStudy (zzr0ck3r):

so A) false, vert line test is to test well if its a function (more technically it is used to make sure a relation is well defined i.e. a=b implies f(a) = f(b)) B) false, the domain of a sequence is \(\mathbb{N}\) the set of natural numbers C) false; surely f(x) = x^2 is a function and we have -2,2 both mapping to 4 D) true: the reason I explained.

OpenStudy (anonymous):

it says that C is also correct... can I argue that?

OpenStudy (zzr0ck3r):

I miss read C. But C still does not give the complete definition of a function. C) False; A function is a relation that is well defined (this is what the option says) AND defined everywhere. This means for all x there is a y such that f(x) = y so A) false, vert line test is to test if its a function (more technically it is used to make sure a relation is well defined i.e. a=b implies f(a) = f(b)) B) false, the domain of a sequence is N the set of natural numbers C) edit above* D) true: the reason I explained.